OpenStudy (anonymous):

integrate (ln x)^6/x

hero (hero):

Lana, maybe I should post my calculus project on here. Maybe then you'll help me

OpenStudy (lalaly):

let u=lnx du=1/x so the integration becomes \[\int\limits{u^6du} = \frac{u^7}{7}\] substitute u=lnx back\[\int\limits{\frac{(lnx)^6}{x}dx} = \frac{(lnx)^7}{7}+C\]

OpenStudy (anonymous):

your awesome!

OpenStudy (anonymous):

what about this one?

OpenStudy (anonymous):

dy/dx=16x(4x^2-1)^3 passing through (1,3)

OpenStudy (lalaly):

whats the question? u want to integrate it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well i have to find the equation

OpenStudy (lalaly):

\[y=\int\limits{16x(4x^2-1)^3dx}\]

OpenStudy (lalaly):

let u= 4x^2-1

OpenStudy (lalaly):

so y becomes\[\frac{1}{2} \times \frac{ (4x^2-1)^5}{5}+C\]

OpenStudy (anonymous):

hmm that isn't an option

OpenStudy (lalaly):

what are the options maybe i mad a lil mistake wait

OpenStudy (lalaly):

the integration is right, u just have t plug in the point that it is passing through

OpenStudy (anonymous):

the options only have powers of 4 and there is no fraction with 4x^2-1

OpenStudy (lalaly):

oh now i know what the mistake is lol

OpenStudy (lalaly):

let u=4x^2-1 du=8xdx \[y=\int\limits{2u^3du}\]\[=\frac{1}{2}\frac{u^4}{4}\]

OpenStudy (lalaly):

i mean\[\frac{1}{2}u^4\]